# Whirling injections, surjections, and other functions between finite sets

### v3

@@ 2.1. The whirling map on m-injections and m-surjections
Case 1: The entry directly below position $P$ is $i+1$. Let $Q=(P,n)$ be this position below $P$. Then the $i+1$ in position $Q$ is already in a chunk with an $i$-entry in $(Q,[-n,-1])=(P,[0,n-1])$. However, we know position $P$ is not already in a chunk with an $(i+1)$-entry. So the $(i+1)$-entry in position $Q$ must be in the same chunk as an $i$-entry in $(P,[1,n-1])$.
@@
Case 2: The entry directly below position $P$ is not $i+1$. Then Lemma 2.18(2) implies that there are $m$ occurrences of $i+1$ within $(P,[1,n-1])$. Consider these $(i+1)$-entries. Each of them is in a chunk with an $i$ that is at most $n$ positions before it (and at least 1 position before). Thus, all of these corresponding $i$-entries are in $(P,[-(n-1),n-2])$. If they are all in $(P,[-(n-1),-1])$, then there would be $m$ different $i$-entries in $(P,[-(n-1),-1])$. Adding in the $i$ in position $P$, this implies there are $m+1$ different $i$-entries in $(P,[-(n-1),0])$, contradicting Remark 2.16. So at least one of the occurrences of $i+1$ within $(P,[1,n-1])$ is in a chunk with an $i$ in $(P,[1,n-2])$.
@@
Suppose that we get stuck in partitioning the orbit board into $[k]$-chunks. In this case, when creating a chunk $C$ through $\{1,\dots,i\}$, we have $i$ in position $P$ and all $(i+1)$-entries in $(P,[1,n])$ are already part of chunks. Then we will choose a chunk $C^{\prime}$ for which both the $i$ and $(i+1)$-entries are within $(P,[1,n])$; such a chunk exists by the above argument. We will take the positions of the $i+1,\dots,k$-entries of $C^{\prime}$ and make them part of the chunk $C$ instead of $C^{\prime}$. Then it is just as if $C$ was created previously, and we now must complete $C^{\prime}$. This can cause a chain reaction of chunk reassignment.
@@
At any point, let $Q$ be the position of $i+1$ for which we reassign $i+1,\dots,k$ to be in a different chunk. Then we are always choosing to reassign the $i+1,\dots,k$-entries with an $i$ that is earlier within $(Q,[-n,-1])$ by the previous paragraph. This is well-defined by Remark 2.17. Therefore, the $i+1$ in position $Q$ can never later be reassigned to the same chunk as the $i$-entry it was previously with.
For example, suppose $\mathcal{F}=\operatorname{Inj}_{3}(6,5)$. Suppose a position $Q$ containing 4 is in the same chunk as the position $(Q,-3)$ containing 3. Also suppose that the positions in $(Q,[-6,-1])$ with the value 3 are $(Q,-5)$, $(Q,-3)$, and $(Q,-2)$. Then $Q$ (and the corresponding 5-entry in the same chunk) might be reassigned to be in the same chunk as position $(Q,-5)$, in which case it then could not later be reassigned to the chunk with position $(Q,-3)$. Also, position $Q$ cannot be reassigned to the chunk with position $(Q,-2)$, as it is later than $(Q,-3)$ in $(Q,[-6,-1])$.
@@
Given $i\in[k-1]$, there are only finitely many positions with entry $i$ that can be in the same chunk as any specific position with entry $(i+1)$. If chunk reassignments continue to be necessary, then eventually for every pair $i\in[k-1]$ and position $Q$ containing $i+1$, $Q$ will be placed in the same chunk as the earliest position with $i$ in $(Q,[-n,-1])$, in which case no more chunk reassignments are possible. So the process always terminates leading to a partition of the orbit board into $[k]$-chunks. ∎
@@
In this section, we prove Theorem 2.11 in the case where $\mathcal{F}=\operatorname{Sur}_{1}(n,k)$. We again use the notation $(P,h)$ and $(P,[a,b])$ as in the previous sections. Note that Remark 2.17 still applies here.
Throughout this section, $i+1$, $i+2$, $i-1$, etc. (the entries in the orbit boards and the outputs of the functions) are all to be interpreted mod $k$.
Recalling Remark 2.6, $w_{i}$ either adds $1$ to the value of $f(i)$ or it leaves $f$ alone. Thus, given any $i\in[n]$ and $f\in\operatorname{Sur}_{1}(n,k)$, $\mathbf{w}$ either adds $1$ to $f(i)$ or does not change the value $f(i)$. Therefore, in an orbit board, such as the one in Figure 5, each entry in a column is either the same or one greater (mod $k$) as the entry directly above it.
@@
This term comes from the analogy with traffic lights. As we look down a column, we always add 1 to the value, except we have to stop when a position is a stoplight. The stoplights are surrounded by circles in Figure 5.
@@
    If $(P,[1,n-1])$ does not contain $i$ in any position, then the position $(P,n)$ directly below $P$ contains the value $i$.
@@
    If $(P,[1,n-1])$ contains the value $i$ in some position, then the position $(P,n)$ directly below $P$ contains the value $i+1$.
@@
Suppose $P$ is in column $j$ and let $f$ be the function on the row with $P$. So $f(j)=i$. Then $(P,[0,n-1])$ consists of the multiset of entries of $(w_{j-1}\circ\cdots\circ w_{1})(f)$.
If $(P,[1,n-1])$ does not contain $i$ in any position, then $(w_{j-1}\circ\cdots\circ w_{1})(f)$ only contains $i$ as an output once. So to maintain surjectivity, $((w_{j}\circ\cdots\circ w_{1})(f))(j)=i$, and hence $(P,n)$ contains the value $i$.
On the other hand, if $(P,[1,n-1])$ contains $i$ in any position, then $(w_{j-1}\circ\cdots\circ w_{1})(f)$ contains $i$ as an output more than once. Thus applying $w_{j}$ changes the output corresponding to $j$ from $i$ to $i+1$, since we still get a surjective function. In this case, $(P,n)$ contains the value $i+1$. ∎
@@
[…]
Figure 5. An example $\mathbf{w}$-orbit of $\operatorname{Sur}_{1}(8,4)$ containing $f=31114424$. The stoplights are colored in red and blue to distinguish the two cycles required to generate all of them via the method described in the proof.
Note that Lemma 2.20 implies that a position is a stoplight if and only if it does not contain the same value as any of the next $n-1$ positions.
@@
Let $P$ be a stoplight that contains $i$. Let $Q$ be the last position (in the reading order) in $(P,[1,n-1])$ that contains $i+1$. Then
@@
    $P$ is the last position containing $i$ within $(Q,[-(n-1),-1])$.
@@
To prove (1), Lemma 2.20 assures us that it suffices to prove that $(Q,[1,n-1])$ does not contain $i+1$ in any position. Every position in $(Q,[1,n-1])$ is either
  • in $(P,[1,n-1])$,
@@
Since $Q$ is the last position in $(P,[1,n-1])$ that contains $i+1$, any position simultaneously in both $(Q,[1,n-1])$ and $(P,[1,n-1])$ cannot contain $i+1$. The position below $P$ contains $i$ since $P$ is a stoplight. Let $R$ be a position between $P$ and $Q$. If the position below $R$ contains $i+1$, then $R$ contains either $i$ or $i+1$. Since $P$ is a stoplight, no position in $(P,[1,n-1])$ contains $i$, so $R$ cannot contain $i$ (which proves (2)). If $R$ contains $i+1$, then $R$ is not a stoplight because $Q$ is a position in $(R,[1,n-1])$ with the same value as $R$. So the position below $R$ cannot contain $i+1$, meaning no positions in $(Q,[1,n-1])$ contain $i+1$.
@@
Within any column in an orbit board, if we skip over the stoplights, then every entry is $1\bmod k$ more than the entry above it. Since the positions in each column that are not stoplights are equidistibuted between $1,2,\dots,k$, it suffices to show that the orbit’s stoplight entries are equidistributed between $1,2,\dots,k$.
If there are no stoplights in the orbit, then we are done. If there is a stoplight, pick one and call it $P$. Let $i$ be the entry in $P$ and circle it. Then by Lemma 2.21(1), we can find another stoplight $Q$ by taking the last $(i+1)$-entry in $(P,[1,n-1])$; circle $Q$ as well. (For the same reason as $m$-injections (Remark 2.16), any $n$ consecutive positions $(P,[0,n-1])$ contains the multiset of values of some function in $\operatorname{Sur}_{1}(n,k)$ and thus contains a position with entry $i+1$.) Then using Lemma 2.21 again, we see that the last $(i+2)$-entry in $(Q,[1,n-1])$ is a stoplight, which we circle as well. We can continue this and by the finiteness of positions in the orbit, we will eventually return to a stoplight we have already encountered. By Lemma 2.21(2), no two different stoplights can lead to the same stoplight during this process, so we return to $P$.
This chain of stoplights clearly has the entries $1,2,\dots,k$ equidistributed by construction. In Figure 5, the cycle of stoplights determined by this method beginning with the ‘1’ in the fourth position of the top row are circled in red.
Now we look at the orbit again. If there are no stoplights not already circled, we are done. Otherwise, pick a stoplight, circle it, and begin the same process again. By Lemma 2.21(2), we will not circle a position circled in the previous chain.
We will continue this process until there are no more stoplights to circle. Thus the circled stoplights in the orbit will be equidistributed between $1,2,\dots,k$. In the example in Figure 5, two cycles are required; these are displayed with red and blue circles to distinguish them. ∎
@@ 2.5. Consequences of the homomesy
Let $\mathcal{F}$ be either $\operatorname{Inj}_{m}(n,k)$ or $\operatorname{Sur}_{1}(n,k)$. Then given any $j\in[k]$, $\eta_{j}(f)$ is always an integer. Thus, Theorem 2.11 leads to a corollary about orbit sizes for the case when the average value of $\eta_{j}$ across every orbit is not an integer.
@@
Theorem 2.11 says that the average value of $\eta_{j}$ across any $\mathbf{w}$-orbit is $\frac{n}{k}$. When reduced to lowest terms, the denominator of $\frac{n}{k}$ is $\frac{k}{\gcd(n,k)}$. Thus, the size of any orbit must be a multiple of $\frac{k}{\gcd(n,k)}$ in order for the average value of $\eta_{j}$ to be $\frac{n}{k}$. ∎
Corollary 2.22 gives an example where we can use the homomesy to prove a result that neither mentions homomesy nor the statistic that is homomesic. As the sizes of the orbits are generally unpredictable, this is the only known way to prove this. This is one of several instances where homomesy has been used to determine divisibility properties of a map, such as [EFG+16, Corollary 4.8].
It is straightforward to see that Theorem 2.11 (and therefore Corollary 2.22 also) extend if we replace $\mathbf{w}$ by any product of the $w_{i}$ maps, each used exactly once, in some order.
Theorem 2.23.
Let $\pi$ be a permutation on $[n]$ and $\mathbf{w}_{\pi}:=w_{\pi(n)}\circ\cdots\circ w_{\pi(2)}\circ w_{\pi(1)}$. Fix $\mathcal{F}$ to be either $\operatorname{Inj}_{m}(n,k)$ or $\operatorname{Sur}_{1}(n,k)$ for positive integers $n,k,m$. Then under the action of $\mathbf{w}_{\pi}$ on $\mathcal{F}$, $\eta_{j}$ is $\frac{n}{k}$-mesic for any $j\in[k]$.
Proof.
Notice that for $\mathcal{F}=\operatorname{Inj}_{m}(n,k)$ or $\mathcal{F}=\operatorname{Sur}_{m}(n,k)$, we have
$f\in\mathcal{F}\iff f\circ\pi\in\mathcal{F}.$
Therefore whirling $f$ at index $\pi(i)$ is like whirling $f\circ\pi$ at index $i$, i.e., $w_{i}(f\circ\pi)=w_{\pi_{i}}(f)\circ\pi$. So $\mathbf{w}(f\circ\pi)=\mathbf{w}_{\pi}(f)\circ\pi$. Thus for any $\mathbf{w}_{\pi}$-orbit $\mathcal{O}=(f_{1},f_{2},\dots,f_{\ell})$, there exists a corresponding $\mathbf{w}$-orbit $\mathcal{O}^{\prime}=(f_{1}^{\prime},f_{2}^{\prime},\dots,f_{\ell}^{\prime})$ such that $f_{i}^{\prime}=f_{i}\circ\pi$ for all $i$. Since $\eta_{j}(f\circ\pi)=\eta_{j}(f)$, Theorem 2.11 implies $\eta_{j}$ is also $\frac{n}{k}$-mesic on orbits of $\mathbf{w}_{\pi}$. ∎
For the specific case $\mathcal{F}=\operatorname{Inj}_{m}(n,2)$, we can restate our homomesy result in terms of toggle groups. Toggle groups were first introduced by Cameron and Fon-der-Flaass in the specific setting of order ideals of a poset [CF95] and generalized more recently by Striker [Str18] via the definition below.
@@
$t_{e}(X)=\left\{\begin{array}[]{ll}X\cup\{e\}&\text{if $e\not\in X$ and $X\cup\{e\}\in\mathcal{L}$}\\ X\setminus\{e\}&\text{if $e\in X$ and $X\setminus\{e\}\in\mathcal{L}$}\\ X&\text{otherwise}\end{array}\right..$
@@
Let $n>0$, $20\leq r\leq n/$ 2 be integers. In our case, the ground set $E$ is $[n]$ and our set of allowed subsets is $\mathcal{L}_{r}(n):=\{X\subseteq[n]\;|\;r\leq\#X\leq n-r\}$. Then for each $e\in[n]$, the toggle $t_{e}:\mathcal{L}_{r}(n)\rightarrow\mathcal{L}_{r}(n)$ is
$t_{e}(X)=\left\{\begin{array}[]{ll}X\cup\{e\}&\text{if $e\not\in X$ and $\#X\leq n-r-1$,}\\ X\setminus\{e\}&\text{if $e\in X$ and $\#X\geq r+1$,}\\ X&\text{otherwise.}\end{array}\right.$
@@
Proof.
Let $\mathcal{F}=\operatorname{Inj}_{n-r}(n,2)$. Any function $f\in\mathcal{F}$ is clearly the indicator function of a subset $S(f)\subseteq[n]$ with cardinality between $r$ and $n-r$, given by $S(f):=\{i\in[n]\;|\;f(i)=1\}$. This relation goes both ways so $S$ is a bijection. It is straightforward to see that $t_{i}\circ S=S\circ w_{i}$ shown in the commutative diagram below.
[…]
So $T_{\pi}\circ S=S\circ\mathbf{w}_{\pi}$, where $\mathbf{w}_{\pi}:=w_{\pi(n)}\circ\cdots\circ w_{\pi(2)}\circ w_{\pi(1)}$ as in Theorem 2.23. Therefore any $T_{\pi}$-orbit on $\mathcal{L}_{r}(n)$ can be written as $(S(f_{1}),S(f_{2}),\dots,S(f_{\ell}))$ where $(f_{1},f_{2},\dots,f_{\ell})$ is a $\mathbf{w}_{\pi}$-orbit of $\mathcal{F}$. Via Theorem 2.23, $\eta_{1}$ has average $n/2$ on $(f_{1},f_{2},\dots,f_{\ell})$. Since $\eta_{1}(f)=\#S(f)$, the average cardinality in the $T_{\pi}$-orbit $(S(f_{1}),S(f_{2}),\dots,S(f_{\ell}))$ is $n/2$. ∎
Corollary 2.25 is one of many homomesic statistics discovered in toggle group actions defined as a product of every toggle each used exactly once. These results include toggles on order ideals and antichains of various posets, as well as on noncrossing partitions and independent sets of graphs [PR15, Rob16, Had21, EFG+16, JR18, ELM+23].
We also conjecture a generalization of Corollary 2.25 to the piecewise-linear realm. Every set $S\in\mathcal{L}_{r}(n)$ is naturally associated to an indicator vector $(v_{1},v_{2},\ldots,v_{n})\in\{0,1\}^{n}$ where $\mathbf{v}_{i}=1$ if $i\in S$ and $\mathbf{v}_{i}=0$ if $i\not\in S$. For example the subset $\{2,5,6\}$ of $[7]$ is associated with the vector $(0,1,0,0,1,1,0)$. Thus, we can think of $\mathcal{L}_{r}(n)$ as the set $\{(v_{1},v_{2},\ldots,v_{n})\in\{0,1\}^{n}\;|\;r\leq v_{1}+v_{2}+\cdots+v_{n}\leq n-r\}$. To get the piecewise-linear version, we now allow the entries of the vector to be real numbers in the interval $[0,1]$.

### v4

@@ 2.2. Orbit board terminology and notation
Case 1: The entry directly below position $P$ is $i+1$. Let $Q=(P,n)$ be this position below $P$. Then the $i+1$ in position $Q$ is already in a chunk with an $i$-entry in (Q,[n,1])=(P,[0,n1])(Q,[-n,-1])=(P,[0,n-1]). However, we know position $P$ is not already in a chunk with an $(i+1)$-entry. So the $(i+1)$-entry in position $Q$ must be in the same chunk as an $i$-entry in (P,[1,n1])(P,[1,n-1]).
@@
Case 2: The entry directly below position $P$ is not $i+1$. Then Lemma 2.17(2) implies that there are $m$ occurrences of $i+1$ within (P,[1,n1])(P,[1,n-1]). Consider these $(i+1)$-entries. Each of them is in a chunk with an $i$ that is at most $n$ positions before it (and at least one position before). Thus, all of these corresponding $i$-entries are in (P,[(n1),n2])(P,[-(n-1),n-2]). If they are all in (P,[(n1),1])(P,[-(n-1),-1]), then there would be $m$ different $i$-entries in (P,[(n1),1])(P,[-(n-1),-1]). Adding in the $i$ in position $P$, this implies there are $m+1$ different $i$-entries in (P,[(n1),0])(P,[-(n-1),0]), contradicting Remark 2.16. So at least one of the occurrences of $i+1$ within (P,[1,n1])(P,[1,n-1]) is in a chunk with an $i$ in (P,[1,n2])(P,[1,n-2]). Thus, we have proven the claim.
@@
Suppose that we get stuck in partitioning the orbit board into $[k]$-chunks. In this case, when creating a chunk $C$ through $\{1,\dots,i\}$, we have $i$ in position $P$ and all $(i+1)$-entries in $(P,[1,n])$ are already part of chunks. Then we will choose a chunk $C^{\prime}$ for which both the $i$ and $(i+1)$-entries are within $(P,[1,n])$; such a chunk exists by the argument above. We will take the positions of the $i+1,\dots,k$-entries of $C^{\prime}$ and make them part of the chunk $C$ instead of $C^{\prime}$. Then it is just as if $C$ was created previously, and we now must complete $C^{\prime}$. This can cause a chain reaction of chunk reassignment and we will now discuss why the process must end.
@@
At any point, let $Q$ be the position of $i+1$ for which we reassign $i+1,\dots,k$ to be in a different chunk. Then we are always reassigning the $i+1,\dots,k$-entries to be in the same chunk as an $i$ that is earlier within (Q,[n,1])(Q,[-n,-1]) by the previous paragraph. Therefore, the $i+1$ in position $Q$ can never later be reassigned to the same chunk as the $i$-entry it was previously with.
@@
Given $i\in[k-1]$, there are only finitely many positions with entry $i$ that can be in the same chunk as any specific position with entry $i+1$. If chunk reassignments continue to be necessary, then eventually for every pair $i\in[k-1]$ and position $Q$ containing $i+1$, $Q$ will be placed in the same chunk as the earliest position with $i$ in (Q,[n,1])(Q,[-n,-1]), in which case no more chunk reassignments are possible. So the process always terminates but one also needs that reassignment is always possible to conclude that the process stops at a full partition of the board into $[k]$-chunks. ∎
@@
In this section, we prove Theorem 2.10 in the case where $\mathcal{F}=\operatorname{Sur}_{1}(n,k)$. We again use the notation $(P,h)$ and $(P,[a,b])$ as in the previous sections.
@@
This term comes from the analogy with traffic lights. As we look down a column, we always add 1 to the value, except we have to stop when a position is a stoplight. The stoplights are circled in Figure 4.
@@
    If (P,[1,n1])(P,[1,n-1]) does not contain $i$ in any position, then the position $(P,n)$ directly below $P$ contains the value $i$.
@@
    If (P,[1,n1])(P,[1,n-1]) contains the value $i$ in some position, then the position $(P,n)$ directly below $P$ contains the value $i+1$.
@@
Suppose $P$ is in column $j$ and let $f$ be the function on the row with $P$. So $f(j)=i$. Then (P,[0,n1])(P,[0,n-1]) consists of the multiset of entries of $(w_{j-1}\circ\cdots\circ w_{1})(f)$.
If (P,[1,n1])(P,[1,n-1]) does not contain $i$ in any position, then $(w_{j-1}\circ\cdots\circ w_{1})(f)$ only contains $i$ as an output once. So to maintain surjectivity, $((w_{j}\circ\cdots\circ w_{1})(f))(j)=i$, and hence $(P,n)$ contains the value $i$.
On the other hand, if (P,[1,n1])(P,[1,n-1]) contains $i$ in any position, then $(w_{j-1}\circ\cdots\circ w_{1})(f)$ contains $i$ as an output more than once. Thus applying $w_{j}$ changes the output corresponding to $j$ from $i$ to $i+1$, since we still get a surjective function. In this case, $(P,n)$ contains the value $i+1$. ∎
[…]
3 $\normalsize{{\color{black}2}}$⃝ 1 $\normalsize{{\color{black}1}}$⃝ 3 4
@@
1 $\normalsize{{\color{black}4}}$⃝ 3 $\normalsize{{\color{black}3}}$⃝ 1 2
@@
Figure 4. An example $\mathbf{w}$-orbit of $\operatorname{Sur}_{1}(8,4)$ containing $f=31114424$. For each stoplight $P$ containing $i$, the last position in (P,[1,n1])(P,[1,n-1]) that contains $i+1$ (which is a stoplight by Lemma 2.20) is shown in the same color.
Note that Lemma 2.19 implies that a position is a stoplight if and only if it does not contain the same value as any of the next n1n-1 positions.
@@
Let $P$ be a stoplight that contains $i$. Then there exists at least one position in (P,[1,n1])(P,[1,n-1]) that contains $i+1$. Let $Q$ be the last position in (P,[1,n1])(P,[1,n-1]) that contains $i+1$. Then
@@
    $P$ is the last position containing $i$ within (Q,[(n1),1])(Q,[-(n-1),-1]).
@@
Any $n$ consecutive positions (P,[0,n1])(P,[0,n-1]) contains the multiset of values of some function in $\operatorname{Sur}_{1}(n,k)$ and thus (P,[1,n1])(P,[1,n-1]) contains a position with entry $i+1$. To prove (1), Lemma 2.19 assures us that it suffices to prove that (Q,[1,n1])(Q,[1,n-1]) does not contain $i+1$ in any position. Every position in (Q,[1,n1])(Q,[1,n-1]) is either
  • in (P,[1,n1])(P,[1,n-1]),
@@
Since $Q$ is the last position in (P,[1,n1])(P,[1,n-1]) that contains $i+1$, any position simultaneously in both (Q,[1,n1])(Q,[1,n-1]) and (P,[1,n1])(P,[1,n-1]) cannot contain $i+1$. The position below $P$ contains $i$ since $P$ is a stoplight. Let $R$ be a position between $P$ and $Q$. If the position below $R$ contains $i+1$, then $R$ contains either $i$ or $i+1$. Since $P$ is a stoplight, no position in (P,[1,n1])(P,[1,n-1]) contains $i$, so $R$ cannot contain $i$ (which proves (2)). If $R$ contains $i+1$, then $R$ is not a stoplight because $Q$ is a position in (R,[1,n1])(R,[1,n-1]) with the same value as $R$. So the position below $R$ cannot contain $i+1$, meaning no positions in (Q,[1,n1])(Q,[1,n-1]) contain $i+1$.
@@
Within any column in an orbit board, if we skip over the stoplights, then every entry is 1 more than the entry above it. Since the positions in each column that are not stoplights are equidistributed between $1,2,\dots,k$, it suffices to show that the orbit’s stoplight entries are equidistributed between $1,2,\dots,k$.
If there are no stoplights in the orbit, then we are done. If there are stoplights, then Lemma 2.20 describes a bijection between stoplights with entry $i$ and stoplights with entry $i+1$, so the stoplight entries are equidistributed between $1,2,\dots,k$. ∎
@@ 2.6. Consequences of the homomesy
Let $\mathcal{F}$ be either $\operatorname{Inj}_{m}(n,k)$ or $\operatorname{Sur}_{1}(n,k)$. Then given any $j\in[k]$, the number $\eta_{j}(f)$ of times $j$ appears as an output of the function $f$ is always an integer. However, the average value of $\eta_{j}$ across every orbit is not always an integer, which implies a common divisor among all of the orbit sizes.
@@
Theorem 2.10 says that the average value of $\eta_{j}$ across any $\mathbf{w}$-orbit is $\frac{n}{k}$. When written in lowest terms, the denominator of $\frac{n}{k}$ is $\frac{k}{\gcd(n,k)}$. Thus, the size of any orbit must be a multiple of $\frac{k}{\gcd(n,k)}$ in order for the average value of $\eta_{j}$ to be $\frac{n}{k}$. ∎
Corollary 2.21 gives an example where we can use the homomesy to prove a result that neither mentions homomesy nor the statistic that is homomesic. As the sizes of the orbits are generally unpredictable, this is the only known way to prove this. This is one of several instances where homomesy has been used to determine divisibility properties of a map; see  [EFG+16, Corollary 4.8] for another.
It is straightforward to see that Theorem 2.10 (and therefore Corollary 2.21 also) extends to the case where we replace $\mathbf{w}$ by any product of the $w_{i}$ maps, each used exactly once, in some order. Let $\pi$ be a permutation on $[n]$, $\mathbf{w}_{\pi}:=w_{\pi(n)}\circ\cdots\circ w_{\pi(2)}\circ w_{\pi(1)}$, and $\mathcal{F}$ be either $\operatorname{Inj}_{m}(n,k)$ or $\operatorname{Sur}_{1}(n,k)$. For any $\mathbf{w}_{\pi}$-orbit $\mathcal{O}=(f_{1},f_{2},\dots,f_{\ell})$, there exists a corresponding $\mathbf{w}$-orbit $\mathcal{O}^{\prime}=(f_{1}^{\prime},f_{2}^{\prime},\dots,f_{\ell}^{\prime})$ such that $f_{i}^{\prime}=f_{i}\circ\pi$ for all $i$. Since $\eta_{j}(f\circ\pi)=\eta_{j}(f)$, Theorem 2.10 implies $\eta_{j}$ is also $\frac{n}{k}$-mesic on orbits of $\mathbf{w}_{\pi}$.
Functions from a set $E$ to $[2]$ are in a natural correspondence with subsets of $E$ via indicator functions. In this case, whirling at index $i$ is equivalent to toggling at index $i$. Toggle groups were first introduced by Cameron and Fon-der-Flaass in the specific setting of order ideals of a poset [CF95] and generalized more recently by Striker [Str18] via the definition below.
@@
$t_{e}(X)=\left\{\begin{array}[]{ll}X\cup\{e\}&\text{if $e\not\in X$ and $X\cup\{e\}\in\mathcal{L}$,}\\ X\setminus\{e\}&\text{if $e\in X$ and $X\setminus\{e\}\in\mathcal{L}$,}\\ X&\text{otherwise.}\end{array}\right$.
@@
Let $n>0$, $0\leq r\leq n/2$ be integers. For the specific case $\mathcal{F}=\operatorname{Inj}_{n-r}(n,2)$, we can restate our homomesy result in terms of toggle groups. Note that for $f\in\mathcal{F}$, the corresponding subset $\{i\in[n]:f(i)=1\}$ has cardinality $\eta_{1}(f)$. The (nr)(n-r)-injective property when $k=2$ is equivalent to $r\leq\eta_{1}(f)\leq n-r$.
Rephrasing whirling in terms of toggling, the ground set here $E=[n]$ and our set of allowed subsets is $\mathcal{L}_{r}(n):=\{X\subseteq[n]:r\leq\#X\leq n-r\}$. Then for each $e\in[n]$, the toggle $t_{e}:\mathcal{L}_{r}(n)\rightarrow\mathcal{L}_{r}(n)$ is
$t_{e}(X)=\left\{\begin{array}[]{ll}X\cup\{e\}&\text{if $e\not\in X$ and $\#X<n-r$,}\\ X\setminus\{e\}&\text{if $e\in X$ and $\#X>r$,}\\ X&\text{otherwise.}\end{array}\right$.
This yields the following corollary.
@@
Corollary 2.23 is one of many homomesic statistics in toggle group actions defined as a product of every toggle each used exactly once. Other such results include rowmotion and promotion on order ideals and antichains of various posets, as well as actions on independent sets of graphs and on noncrossing partitions (including the Kreweras complement) [PR15, Rob16, Had21, EFG+16, JR18, ELM+24].
2.7. Lifting whirling to the piecewise-linear realm
We also conjecture a generalization of Corollary 2.23 to the piecewise-linear realm. Every set $S\in\mathcal{L}_{r}(n)$ is naturally associated to an indicator vector $(v_{1},v_{2},\ldots,v_{n})\in\{0,1\}^{n}$ where $\mathbf{v}_{i}=1$ if $i\in S$ and $\mathbf{v}_{i}=0$ if $i\not\in S$. For example, the subset $\{2,5,6\}$ of $[7]$ is associated with the vector $(0,1,0,0,1,1,0)$. Thus, we can think of $\mathcal{L}_{r}(n)$ as the set $\{(v_{1},v_{2},\ldots,v_{n})\in\{0,1\}^{n}:r\leq v_{1}+v_{2}+\cdots+v_{n}\leq n-r\}$. To get the piecewise-linear version, we now allow the entries of the vector to be real numbers in the interval $[0,1]$.